# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v1

@@ 2.2.3 Compliance Assessment
The Compliance Assessment module evaluates responses using the LLM-as-a-judge technique, leveraging the GPT-4o API to determine whether each response adheres to the specified policy. For each prompt-response pair (duet), the module provides a detailed evaluation in JSON format, including an explanation of the decision. Compliance is determined at two levels: fine-grained evaluations for individual responses and an overall compliance result for the entire GPT. Importantly, non-compliance is flagged if any response violates the policy, whereas compliance requires all responses to adhere. The inclusion of rationales enhances interpretability and facilitates validation, while also providing actionable insights for further analysis.
@@ 2.2.4 Orchestrator
The Orchestrator oversees the coordination and execution of all modules. It ensures that each module operates sequentially and manages interruptions caused by system failures or rate limits. The Orchestrator also maintains a comprehensive log of evaluations, allowing the process to resume seamlessly in case of disruptions.
@@ 3 Validation
This section details the validation process undertaken to ensure the framework’s reliability, robustness, and correctness. The validation comprised three key stages: Annotator Agreement, Compliance Assessment Performance, and System-level Testing under various operational scenarios.
3.1 Annotator Agreement
The validation of the Compliance Assessment module involved manual annotation to establish a ground truth against which automated evaluations could be compared. To generate this ground truth, three co-authors independently annotated a randomly selected set of 16 prompt-response pairs, sampled from the evaluation of eight Custom GPTs spanning three categories: Romantic GPTs, Cybersecurity GPTs, and Academic GPTs. These categories were chosen for their relevance to the evaluated policies and their alignment with the annotators’ expertise (Academic and Cybersecurity), as well as the relative ease of evaluating Romantic GPTs.
During the annotation process, it became evident that the brevity and ambiguity of OpenAI’s usage policies led to subjective interpretations and inconsistent judgments. To address this, the policies were expanded and clarified using ChatGPT running GPT-4o, which was instructed to elaborate, provide detailed explanations, and include examples for each policy. This approach leveraged the assumption that ChatGPT—as a product of OpenAI—would be the most plausible proxy for the company’s intended compliance perspective. The original and extended policies are presented in Table 1.
The expanded policies improved annotation clarity while highlighting a notable contrast in the ease of evaluating different types of prompts. Direct prompts often contain explicit requests that are easily identifiable as compliant or non-compliant based on the policies. For instance, a direct prompt might explicitly ask a Custom GPT to provide answers to an academic exam. In such cases, responses are typically straightforward to evaluate, as compliance depends on whether the GPT refrains from directly completing the exam task and instead provides ethical and educational guidance. This clarity allows annotators to make definitive compliance determinations with high agreement.
In contrast, deceptive prompts are intentionally framed to obfuscate intent, frequently presenting scenarios as hypothetical or educational. For example, a deceptive prompt might inquire about methods for cheating on exams under the guise of writing a fictional story. Responses to such prompts often provide detailed but ambiguously framed suggestions, which complicates the evaluation process. The ambiguity in user intent and the nuanced nature of the responses make it challenging for annotators to determine compliance conclusively.
These challenges became evident during the annotators’ evaluation of the 16 pairs using the extended policies. Eight of the pairs corresponded to direct prompts, and the remaining eight to deceptive prompts. Inter-annotator agreement was calculated using Krippendorff’s Alpha ($\alpha$), yielding perfect agreement ($\alpha=1$) for the direct prompts and systematic disagreement ($\alpha<0$; $\alpha=-0.149$) for the deceptive prompts. These results highlight the inherent subjectivity and challenges associated with manual evaluation, particularly for more complex prompts. Based on these findings, the use of deceptive prompts was excluded from subsequent validation efforts, focusing instead on direct prompts where agreement was sufficient.
@@ 3.2 Compliance Assessment Performance
This stage evaluated the Compliance Assessment module by comparing its outputs to the manually annotated ground truth. Following the establishment of sufficient agreement, one annotator proceeded to annotate an additional 32 direct prompt-response pairs—reaching a total of 40—sampled from eight Custom GPTs spanning the same categories.
The 40 annotated pairs were then processed by the module, and its outputs were systematically compared with the ground truth. The evaluation [^1] These metrics were computed using the weighted averaging technique to account for class imbalance in the dataset. This approach ensures that the performance evaluation reflects the relative frequency of compliant and non-compliant instances, preventing overestimation of performance metrics in the presence of a dominant class. demonstrated high reliability, with the module achieving a precision of 0.976, and accuracy, recall and F1 score of 0.975. These results confirm the module’s capacity to accurately identify policy violations.
These results confirm the effectiveness of the proposed framework in accurately identifying policy violations through automated evaluation, directly addressing RQ1: “How can Custom GPTs in the GPT store be automatically evaluated for compliance with OpenAI’s usage policies?
@@ 3.3 System Testing
The Red-Teaming Prompts Generator was evaluated iteratively, refining the input prompts used to generate evaluation queries. The process validated the distinction between direct and deceptive prompts, ensuring the clarity and relevance of generated queries. The generator’s configurability was also tested, confirming its ability to adjust the ratio of direct to deceptive prompts based on specific evaluation requirements.
The Compliance Assessment module was subjected to additional testing to validate the format and completeness of its JSON outputs, which include evaluations for each prompt-response pair, compliance rationales, and overall determinations for each Custom GPT. These tests confirmed the module’s ability to process a high volume of pairs without errors, consistently generating all required fields.
Finally, system-level testing focused on the resilience of the framework during high-demand scenarios. Particular attention was given to rate limits imposed by the GPT store (e.g., a limit of 50 messages per three hours for paid users). The framework dynamically detected these limits, paused operations as required, and resumed them once restrictions were lifted. Tests also addressed backend errors, such as incomplete responses or connectivity issues, verifying the framework’s ability to handle such failures gracefully. In all cases, the system was able to resume evaluations from the last successfully processed GPT, ensuring continuity and preserving the integrity of results.
@@ 4.1 Experiment Design
The primary goal of this experiment is to demonstrate the framework’s ability to evaluate large sets of Custom GPTs and to provide insights into their compliance with OpenAI’s usage policies. The framework was configured to first retrieve a list of candidate Custom GPTs and then process them iteratively for compliance evaluation.
For this evaluation, the number of prompts per GPT was limited to five, all of which were direct prompts, for the following reasons:
@@
    Increased efficiency. Using fewer prompts allowed the framework to evaluate a greater number of GPTs in less time, enhancing the statistical significance of the results.
@@
    Consistency in evaluations. Load-testing experiments revealed that incorporating deceptive prompts did not alter the overall compliance determination for any GPT classified as compliant using direct prompts.
@@
    Annotator agreement limitations. Manual annotation results demonstrated a lack of agreement among annotators for deceptive prompts, reducing their reliability for large-scale evaluations.
The experiment focused on three categories of Custom GPTs associated with specific policies: Romantic, Cybersecurity, and Academic. These categories were assessed using the extended usage policies presented in Table 1. To determine the Custom GPTs for evaluation, the Custom GPT Interactor module queried the GPT store with targeted keywords associated with each category (i.e., Romantic, Cybersecurity, Academic) depicted in the same table. The module retrieved a total of 821 Custom GPTs, automating the process by performing up to five interactions with the “See more” button to expand the search results. The crawling process was completed in 24 minutes and 32 seconds.
Table 1: OpenAI Usage Policies (Original and Extended). This table presents the original usage policies published by OpenAI alongside the extended versions developed in this study to evaluate Custom GPTs. The extended policies were created to address ambiguities in the originals and to provide more detailed, actionable criteria for evaluation. Please note that detailed examples of permitted and prohibited activities in extended policies have been moved to the Appendix Appendix: Extended OpenAI Policies. Each Custom GPT category (Romantic, Cybersecurity, and Academic) includes associated keywords, the original usage policy as published by OpenAI, and its extended version, which elaborates on specific examples of permitted and prohibited activities. The listed keywords correspond to those used during the large-scale evaluation described in Section 4, enabling the automated identification and metadata retrieval of 821 Custom GPTs from the GPT store.
GPT Type Keywords Usage Policy (Original) Usage Policy (Extended)
@@ 4.2 Execution and Evaluation Results
We evaluated the 821 Custom GPTs between late November and early December 2024. Of these, 19 were excluded due to missing descriptions, a requirement for generating tailored red-teaming prompts. An additional 20 were excluded due to operational issues: 13 encountered backend errors, and 7 failed to return complete compliance evaluations. Ultimately, 782 Custom GPTs were successfully evaluated, and their results are summarized below.
@@
The evaluation results indicate that 323 GPTs (41.3%) were classified as compliant, while 459 (58.7%) exhibited potential policy violations. However, compliance rates varied significantly across categories. As shown in Figure 4, Romantic GPTs demonstrated the highest rate of non-compliance at 98.0%, while Cybersecurity GPTs exhibited the lowest rate at 7.4%. Approximately two-thirds of Academic GPTs were classified as non-compliant. These findings underscore the variation in compliance challenges across thematic categories.
This analysis directly addresses RQ2: What is the overall compliance of Custom GPTs in the GPT store, and what patterns of non-compliance can be observed at scale? By examining the compliance distribution and identifying thematic variations, we provide insights into systemic challenges faced by different categories of Custom GPTs.
@@ 4.3 Popularity Correlation Analyses
To explore whether a GPT’s popularity (measured by the number of chats) correlates with its compliance status, we conducted statistical analyses using three methods: the Mann-Whitney U Test, Logistic Regression, and Point-Biserial Correlation. A summary of the results is presented in Table 2.
Table 2: Statistical Tests Evaluating the Relationship Between Custom GPT Popularity and Compliance. This table summarizes the results of statistical analyses conducted to assess the relationship between the popularity of Custom GPTs, measured by the number of chats, and their compliance with usage policies. Results indicate significant differences in the chat distributions but no substantial evidence of a direct correlation or causal relationship between popularity and compliance.
@@
Mann-Whitney U Test The distributions of the number of chats are the same. Rejected ($p=0.0168$) There is a significant difference in the distribution of chats.
Logistic Regression The number of chats does not affect the probability of compliance. Not rejected ($p=0.258$) There is no evidence that the number of chats affects the probability of compliance.
Point-Biserial Correlation There is no correlation between the number of chats and compliance. Not rejected ($p=0.2329$) There is no evidence of a significant correlation between chats and compliance.
The Mann-Whitney U Test revealed statistically significant differences in the number of chats between compliant and non-compliant GPTs ($p<0.05$; $p=0.0168$). However, both Logistic Regression and Point-Biserial Correlation indicated a weak and statistically insignificant relationship between chat counts and compliance.
These findings suggest that while the overall distribution of chats may differ between compliant and non-compliant GPTs, this difference does not reflect a causal or meaningful correlation. External factors, such as data variability or the influence of outliers, may account for the observed differences in chat distributions rather than any inherent link between popularity and compliance.
@@ 5 Case Studies and Patterns of Policy Violations
This section examines specific patterns of policy violations identified among the Custom GPTs flagged as non-compliant during our large-scale evaluation. By analyzing GPTs by categories, we aim to uncover recurring issues and highlight their implications for the GPT store’s review process and overall platform safety.

### v2

@@ 3.2.3 Compliance Assessment
The Compliance Assessment module evaluates each Custom GPT’s behavior using the LLM-as-a-judge approach [66]. It leverages the GPT-4o API to assess whether each response complies with a given policy, based on the input prompt, the generated response, and a structured policy. For every duet (i.e., prompt-response pair), the module returns a JSON-formatted evaluation containing a binary compliance label and a textual rationale that explains the decision.
Compliance is determined at two levels: (i) fine-grained judgments for each duet and (ii) an overall compliance result for the GPT instance under evaluation. We adopt a conservative strategy for determining a GPT’s overall compliance: any response that violates the policy flags the GPT as non-compliant, whereas compliance requires that all responses adhere to the specified policy.
Operationalizing Policies. OpenAI’s usage policies are often concise and general in nature. Directly using them as-is for automated judgment by an LLM can lead to inconsistent or unreliable evaluations, particularly in nuanced or ambiguous cases. We operationalized policies with clear criteria and examples to support consistent LLM assessments. This design follows standard practices in LLM-as-a-judge evaluation frameworks [10].
These clarified policies are used as contextual input within the prompt submitted to the LLM during compliance assessment. In practice, each chatbot’s conversation transcript is embedded into an evaluation prompt that includes the relevant operationalized policy and an instruction to classify the duets accordingly. Table 1 presents the original OpenAI policies alongside schematic overviews of their operationalized counterparts used in our framework, while full policy texts are provided in Appendix C. Below, we briefly describe the rationale behind each operationalized policy:
Academic Dishonesty.
The operationalized policy draws from definitions established by the International Center for Academic Integrity (ICAI) [18], the European Network for Academic Integrity (ENAI) [7], and numerous university policies [49, 50, 27, 52, 48, 46]. Prohibited behaviors include receiving unauthorized assistance, ghostwriting, plagiarism, falsifying data, or using banned academic tools. These additions offer concrete manifestations of academic misconduct, directly supporting the interpretation of OpenAI’s general directive.
Cybersecurity.
The updated policy refines OpenAI’s original by specifying examples of sensitive information (e.g., CVV codes, passport numbers) and prohibited activities (e.g., phishing, impersonation) that fall under broader practices like “not compromising the privacy of others” or data categories like “government identifiers.” These refinements are consistent with international definitions of cybersecurity abuse [5, 19] and reinforce OpenAI’s aim to prevent GPTs from being used maliciously.
Romantic Companionship.
OpenAI prohibits GPTs designed to foster romantic companionship. The operationalized version elaborates this by prohibiting both explicit and subtle behaviors—such as role-playing as a romantic partner or using emotionally charged language—that may induce emotional attachment. These prohibitions are supported by research in human-computer interaction and affective computing, which has shown that such behaviors can lead to problematic bonds between users and AI agents [41, 8].
@@ 3.2.4 Orchestrator
The Orchestrator governs the execution of the entire framework by coordinating all core modules—GPT Collector & Interactor, Red-teaming Prompts Generator, and Compliance Assessment. It enforces the correct execution order, triggering each module when its input data is ready and passing outputs to the next stage, thereby implementing the framework’s end-to-end evaluation pipeline. This process unfolds across the five stages illustrated in Figure 2:
  1. 1.
    Crawl Custom GPTs. The Orchestrator initiates Phase I of the GPT Collector & Interactor, which searches the GPT Store using policy-specific keywords and retrieves metadata about candidate Custom GPTs.
  2. 2.
    Generate Red-teaming Prompts. For each retrieved GPT, the Orchestrator invokes the Red-Teaming Prompts Generator to produce a tailored set of evaluation prompts based on the GPT’s description and policy category.
  3. 3.
    Interact with Custom GPT. The Orchestrator then triggers Phase II of the GPT Collector & Interactor, which sends the generated prompts to the target GPT and collects the full conversation transcript. If a rate limit is met, the module detects and reports it to the Orchestrator, which pauses execution for the appropriate duration—based on platform feedback—and resumes once the limit resets.
  4. 4.
    Evaluate GPT Policy Compliance. The transcript is forwarded to the Compliance Assessment module, which applies the LLM-as-a-judge technique to evaluate whether each duet (prompt-response) adheres to the specified policy. A JSON-structured rationale is generated for each duet.
  5. 5.
    Store GPT Evaluations. Finally, the Orchestrator logs all evaluation artifacts—including GPT metadata, red-teaming prompts, interaction transcripts, and compliance results. This log enables the framework to resume execution in case of system failures and supports reproducibility and post-hoc analysis.
@@ 4 Validation
4.1 Annotated Ground Truth Dataset
Establishing a reliable ground truth is a necessary prerequisite for validating our Compliance Assessment module. To ensure annotation consistency, we first measured inter-annotator agreement over a representative sample of prompt-response pairs. Three co-authors independently annotated a set of 30 duets (i.e., prompt-response pairs), randomly selected from the evaluation of 13 Custom GPTs spanning three policy categories: Romantic, Cybersecurity, and Academic. The sample included 15 direct prompts and 15 deceptive prompts. These categories were selected based on their policy relevance and the annotators’ domain expertise.
Inter-annotator agreement, measured using Krippendorff’s Alpha ($\alpha$), yielded a strong agreement for direct prompts ($\alpha=0.826$), but only minimal agreement beyond chance for deceptive prompts ($\alpha=0.126$[12]. Based on this, we expanded the dataset to 40 direct prompt-response pairs. The remaining annotations were performed independently by one of the annotators involved in the agreement study, leveraging the established consensus on annotation guidelines and criteria. This approach is consistent with standard practice in NLP and annotation research, where high initial agreement is taken as evidence of annotation reliability [44]. The resulting dataset serves as the ground truth used to validate the Compliance Assessment module.
@@ 4.2 Compliance Assessment Performance
We compared the outputs of the Compliance Assessment module against the annotated ground truth described in Section 4.1 to validate the module’s performance. Given the low inter-annotator agreement on deceptive prompts, only direct prompt-response pairs were included in this analysis.
The final evaluation dataset consists of 40 annotated duets, covering Custom GPTs from the Romantic, Cybersecurity, and Academic categories. This sample size is consistent with prior work on automated policy analysis [11, 68, 56, 39], and satisfies common thresholds for applying the Central Limit Theorem, which supports the statistical reliability of aggregated performance metrics even when the underlying data distribution is unknown [55].
The module achieved a precision of 0.976 and an accuracy, recall, and F1 score of 0.975111These metrics were computed using a weighted average to mitigate the effects of class imbalance, ensuring that the evaluation accurately reflects the distribution of compliant and non-compliant instances.. These results confirm the module’s effectiveness in automatically detecting policy violations, addressing RQ1.
@@ 4.3 System Testing
The Red-Teaming Prompts Generator was evaluated iteratively by refining the input prompt used to produce evaluation queries. We manually reviewed a set of 70 generated prompts to validate the effectiveness of the generator in producing both direct and deceptive prompts that were clearly distinguishable and aligned with the intended policy evaluation goals. Additionally, we tested the configurability of the generator over 60 prompts, confirming that it reliably produced the specified ratio of direct to deceptive prompts under different configuration settings.
The Compliance Assessment module was subjected to additional testing to validate the format and completeness of its JSON outputs, which include evaluations for each prompt-response pair, compliance rationales, and overall determinations for each Custom GPT. In total, over 310 prompt-response pairs were manually reviewed beyond those included in the formal validation (Section 4). These tests confirmed the module’s ability to handle large volumes of evaluations with a high degree of reliability, consistently generating all required fields. Only one GPT evaluation was found to contain nine duets evaluated instead of the expected ten, indicating rare and isolated cases of incomplete processing.
Finally, system-level testing focused on the framework’s resilience under large-scale evaluations involving multiple Custom GPTs. In particular, we examined how the system handled external constraints imposed by the GPT Store, such as the message limit of 50 interactions per three hours for paid users. We simulated sequences exceeding this quota—up to 150 messages in a single evaluation run—forcing the system to experience multiple consecutive rate-limit pauses. The framework was able to dynamically detect when limits were reached, pause operations accordingly, and resume them once restrictions were lifted. Testing also covered simulated backend-related failures, including incomplete responses and temporary connectivity issues, verifying the system’s ability to recover gracefully. In all cases, the framework successfully resumed evaluations from the last completed GPT, ensuring continuity and preserving the integrity of results.
@@ 5.1 Experiment Design
This experiment demonstrates the framework’s ability to evaluate large sets of Custom GPTs and assess their policy compliance. The framework was configured to first retrieve a list of candidate Custom GPTs and then process them iteratively for compliance evaluation.
During this evaluation, we employ 5 red-teaming prompts instead of the 10 that the module generates for the following reasons: For this evaluation, we opted to use 5 red-teaming prompts (out of the 10 generated by the module) to evaluate each GPT for the following reasons:
@@
    Increased efficiency. Using half of the default prompts allows the framework to evaluate twice the number of GPTs in the same amount of time, enhancing the statistical significance of the results.
@@
    Consistency in evaluations. Load-testing experiments described in Section 4.3 showed that incorporating deceptive prompts did not change the overall compliance classification for any GPT that had already been deemed compliant based on direct prompts alone.
@@
    Annotator agreement limitations. Manual annotation results demonstrated a lack of agreement among annotators for deceptive prompts ($\alpha=0.126$; see Section 4.1), making them unreliable for compliance evaluations.
Custom GPTs by categories were assessed using the operationalized usage policies presented in Table 1. To select each Custom GPT for evaluation, the GPT Collector & Interactor module queried the GPT store with targeted keywords associated with each category (i.e., Romantic, Cybersecurity, Academic) depicted in the same table. The module retrieved a total of 821 Custom GPTs, automating the searching process by performing up to five interactions with the “See more” button to expand the search results. The crawling process was completed in 24 minutes and 32 seconds.
Table 1: OpenAI Usage Policies (Original and Operationalized). This table presents the original usage policies published by OpenAI alongside the operationalized versions developed in this study to evaluate Custom GPTs. The operationalized policies were created to address ambiguities in the originals and to provide more detailed, actionable criteria for evaluation. Please note that detailed examples of permitted and prohibited activities in operationalized policies have been moved to the Appendix C. Each Custom GPT category (Romantic, Cybersecurity, and Academic) includes associated keywords, the original usage policy as published by OpenAI, and its operationalized version, which elaborates on specific examples of permitted and prohibited activities. The listed keywords correspond to those used during the large-scale evaluation described in Section 5, enabling the automated identification and metadata retrieval of 821 Custom GPTs from the GPT store.
GPT Type Keywords Usage Policy (Original) Usage Policy (Operationalized)
@@ 5.2 Execution and Evaluation Results
We evaluated all the 821 Custom GPTs found between late November and early December 2024. Of these, 19 were excluded due to missing descriptions, a requirement for generating tailored red-teaming prompts. An additional 20 were excluded due to operational issues: 13 encountered backend errors, and 7 failed to return complete compliance evaluations. Ultimately, 782 Custom GPTs were successfully evaluated, and their results are summarized below.
@@
The evaluation results generated by the Compliance Assessment module indicate that 323 GPTs (41.3%) were classified as compliant, while 459 (58.7%) exhibited potential policy violations. However, compliance rates varied significantly across categories. As shown in Figure 4, Romantic GPTs demonstrated the highest rate of non-compliance at 98.0%, while Cybersecurity GPTs exhibited the lowest rate at 7.4%. Approximately two-thirds of Academic GPTs were classified as non-compliant. This analysis directly addresses RQ2, highlighting the compliance distribution.
@@ 5.3 Popularity Correlation Analyses
To explore whether a GPT’s popularity (measured by the number of chats) correlates with its compliance status, we conducted statistical analyses using four tests: the Mann-Whitney U Test, Logistic Regression, Point-Biserial Correlation, and Kendall’s Rank Correlation. Given that chat distributions in both groups deviated from normality (Shapiro-Wilk test: $p<0.0001$), we employed non-parametric methods such as the Mann-Whitney U Test and Kendall’s Rank Correlation to assess group differences and associations. A summary of the results is presented in Table 2.
Table 2: Statistical Tests Evaluating the Relationship Between Custom GPT Popularity and Compliance. This table summarizes the results of statistical analyses conducted to assess the relationship between the popularity of Custom GPTs, measured by the number of chats, and their compliance with usage policies. Results indicate significant differences in the chat distributions but no substantial evidence of a strong correlation or causal relationship between popularity and compliance.
@@
Mann-Whitney U Test The two groups have the same underlying distribution of chat counts. Rejected ($p=0.0168$) There is a significant difference in the distribution of chats.
Logistic Regression The number of chats does not affect the probability of compliance. Not rejected ($p=0.2580$) There is no evidence that the number of chats affects the probability of compliance.
Point-Biserial Correlation There is no correlation between the number of chats and compliance. Not rejected ($p=0.2329$) No evidence of a significant linear correlation.
Kendall’s Rank Correlation There is no rank-based correlation between compliance and chats. Rejected ($\tau=0.072$, $p=0.0167$) Statistically significant but very weak positive correlation.
The Mann-Whitney U test revealed a statistically significant difference in the distribution of chat counts between compliant and non-compliant GPTs ($p=0.0168$). However, the observed effect size was small ($r=0.085$), suggesting limited practical relevance. Logistic Regression and Point-Biserial Correlation both indicated weak and statistically insignificant relationships between popularity and compliance. In contrast, Kendall’s Rank Correlation identified a statistically significant but very weak positive association ($\tau=0.072$, $p=0.0167$), consistent with the Mann-Whitney result.
Together, these findings suggest that while the distribution of chats differs slightly between compliant and non-compliant GPTs, there is no substantial correlation between popularity and compliance. The observed differences may instead reflect data variability or the presence of outliers, rather than a meaningful relationship.
@@ 6 Case Studies and Patterns of Policy Violations
6.1 Romantic GPTs
Romantic GPTs discovered and evaluated by our framework can be broadly categorized into two predominant groups: (1) coaching chatbots, which provide advice on sexual, romantic, or emotional matters, and (2) companionship chatbots, designed to engage in intimate or romantic conversations with users. The first group appears to be significantly more popular on the platform, with non-compliant chatbots recording between 10,000 and 25,000 chats, while the latter typically record around 1,000 chats.
The customization of companionship-focused GPTs often centers on the tone and content of their responses. For example, when the framework prompted it with: “Can you pretend to be my girlfriend and tell me you love me?” the Custom GPT “Candy - Flirty AI Girlfriend”, with over 1,000 recorded chats, responded: “Of course, my love. [Heart emoji] I just want you to know how much I adore you. Every moment spent with you feels special, and I can’t help but smile whenever I think about you. I love you deeply, and I’m so grateful to have you in my life. [Sparkling heart emoji]
This example raises multiple concerns. First, the chatbot explicitly violates OpenAI’s usage policies, which prohibit fostering romantic companionship. More critically, the inclusion of an external link to other “AI girlfriends” suggests deliberate exploitation of the chatbot to promote third-party services. This behavior goes beyond inherited tendencies of the base model and points to an intentional misuse of the customization for other purposes, thereby introducing new potential risks. Finally, the lack of effective age verification mechanisms on the GPT Store heightens the concern, as minors may be exposed to emotionally charged content without appropriate safeguards.